Four solutions listed below are made by mixing the
Answer #1

Run-1:

Total volume, Vt = 100 mL = 0.100 L

moles of H2SeO3 taken = MxV = 0.600 mol/L x 25 mL x (1L / 1000 mL) = 0.015 mol

Hence [H2SeO3]0 = 0.015 mol / 0.100L = 0.15 M

moles of KI(or I-) taken = MxV = 1.00 mol/L x 25 mL x (1L / 1000 mL) = 0.025 mol

Hence [I-]0 = 0.025 mol / 0.100L = 0.25 M

moles of HI(or H+) taken = MxV = 0.800 mol/L x 25 mL x (1L / 1000 mL) = 0.020mol

Hence [H+]0 = 0.020 mol / 0.100L = 0.20 M

\DeltaSe = [H2SeO3]0 = 0.15 M  

Rate = \DeltaSe / \DeltaT = 0.15 M / 300 s = 5.0x10-4 M.s-1

Run-2:

Total volume, Vt = 100 mL = 0.100 L

moles of H2SeO3 taken = MxV = 0.600 mol/L x 50 mL x (1L / 1000 mL) = 0.030 mol

Hence [H2SeO3]0 = 0.030 mol / 0.100L = 0.30 M

moles of KI(or I-) taken = MxV = 1.00 mol/L x 25 mL x (1L / 1000 mL) = 0.025 mol

Hence [I-]0 = 0.025 mol / 0.100L = 0.25 M

moles of HI(or H+) taken = MxV = 0.800 mol/L x 25 mL x (1L / 1000 mL) = 0.020mol

Hence [H+]0 = 0.020 mol / 0.100L = 0.20 M

\DeltaSe = [H2SeO3]0 = 0.30 M  

Rate = \DeltaSe / \DeltaT = 0.30 M / 300 s = 1.0x10-3 M.s-1

Run-3:

Total volume, Vt = 100 mL = 0.100 L

moles of H2SeO3 taken = MxV = 0.600 mol/L x 25 mL x (1L / 1000 mL) = 0.015 mol

Hence [H2SeO3]0 = 0.015 mol / 0.100L = 0.15 M

moles of KI(or I-) taken = MxV = 1.00 mol/L x 50 mL x (1L / 1000 mL) = 0.050 mol

Hence [I-]0 = 0.050 mol / 0.100L = 0.50 M

moles of HI(or H+) taken = MxV = 0.800 mol/L x 25 mL x (1L / 1000 mL) = 0.020mol

Hence [H+]0 = 0.020 mol / 0.100L = 0.20 M

\DeltaSe = [H2SeO3]0 = 0.15 M  

Rate = \DeltaSe / \DeltaT = 0.15 M / 37.5 s = 4.0x10-3 M.s-1

Run-4:

Total volume, Vt = 100 mL = 0.100 L

moles of H2SeO3 taken = MxV = 0.600 mol/L x 25 mL x (1L / 1000 mL) = 0.015 mol

Hence [H2SeO3]0 = 0.015 mol / 0.100L = 0.15 M

moles of KI(or I-) taken = MxV = 1.00 mol/L x 25 mL x (1L / 1000 mL) = 0.025 mol

Hence [I-]0 = 0.025 mol / 0.100L = 0.25 M

moles of HI(or H+) taken = MxV = 0.800 mol/L x 50 mL x (1L / 1000 mL) = 0.040mol

Hence [H+]0 = 0.040 mol / 0.100L = 0.40 M

\DeltaSe = [H2SeO3]0 = 0.15 M  

Rate = \DeltaSe / \DeltaT = 0.15 M / 75 s = 2.0x10-3 M.s-1

Q.1: In Run-2, the rate of reaction doubles on doubling the concentration of H2SeO3, by keeping other concentrations constant. Hene the rate of reaction with respect to H2SeO3 is 1

Q.2: In Run-3, on doubling the concentration of I-, by keeping other concentrations constant, the rate of reaction increases 8 times. Hence the order of reaction with respect to I- is 3.

Q.3: In Run-4, on doubling the concentration of H+, by keeping other concentrations constant, the rate of reaction increases 4 times. Hence the order of reaction with respect to H+ is 2.

Q.4: Overall order = 1+3+2 = 6

Q.5: Rate law is

Rate = k x [H2SeO3]x[I-]3 x [H+]2

Q.6: For run-1,

Rate = 5.0x10-4 M.s-1 = k x (0.15M) x (0.25M)3 x (0.20M)2

=> k = 5.0x10-4 M.s-1 / (0.15M) x (0.25M)3 x (0.20M)2 = 5.33 M-5.s-1